# Mind the (DH) Gap! A Contrast in Risky Choices Between Reasoning and Conversational LLMs

### v1

@@ Abstract
The use of large language models either as decision support systems, or in agentic workflows, is rapidly transforming the digital ecosystem. However, the understanding of LLM decision-making under uncertainty remains limited. We initiate a comparative study of LLM risky choices along two dimensions: (1) prospect representation (explicit vs. experience based) and (2) decision rationale (explanation). Our study, which involves 20 frontier and open LLMs, is complemented by a matched human subjects experiment, which provides one reference point, while an expected payoff maximizing rational agent model provides another. We find that LLMs cluster into two categories: reasoning models (RMs) and conversational models (CMs). RMs tend towards rational behavior, are insensitive to the order of prospects, gain/loss framing, and explanations, and behave similarly whether prospects are explicit or presented via experience history. CMs are significantly less rational, slightly more human-like, sensitive to prospect ordering, framing, and explanation, and exhibit a large description-history gap. Paired comparisons of open LLMs suggest that a key factor differentiating RMs and CMs is training for mathematical reasoning.
@@ 1 Introduction
However, LLMs are neither explicitly designed to maximize expected payoff, nor necessarily behave like humans. Rather, their competence is an emergent property of scale and training paradigms including next-word prediction, instruction fine-tuning, human preference alignment, and mathematical reasoning (Du et al., 2024; Guo et al., 2025; Ouyang et al., 2022). The net effect of this soup of training methods on LLM economic decision-making under uncertainty remains unclear. In response, a literature has begun to emerge that aims to investigate LLM decision-making, particularly under uncertainty. However, there remain conflicting accounts, for example, of the extent LLMs are rational or align with prospect theory and human behavior (Horton, 2023; Jia et al., 2024; Ross et al., 2024; Wang et al., 2025; Payne, 2025). More fundamentally, given the broad variety of training paradigms, goals, and architectures, it is unclear whether a straightforward account of LLM economic decision making is even possible.
@@
We study a suite of 20 frontier and open LLMs, along with human subjects’ choices over the same prospects and treatments. LLMs are analyzed vis-a-vis two references: 1) the human subjects pool, and 2) the idealized economicus rational agent. Within the scope of our experiments, our key findings are:
@@
    All models exhibit an interface analogue of the DE gap (DH gap). This gap is considerable for CMs, but modest for RMs.
@@
Related Work. LLMs are increasingly studied as objects of behavioral analysis Horton (2023); Dillion et al. (2023). This line of “machine psychology” research emphasizes controlled perturbations and cautions against importing human psychological constructs without careful study Hagendorff et al. (2023); Ivanova (2023). Jia et al. (2024) analyze LLM decisions by eliciting choices and fitting prospect-theoretic models; subsequent work extends this by examining epistemic markers Wang et al. (2025), application contexts Payne (2025), and persona matching  Liu et al. (2025). Horowitz and Plonsky (2025) study a description-experience gap in LLMs through repeated choice, unlike the passive sampling used in our setting. Apart from the bounded rationality focus, Chen et al. (2023) instead focuses on rationality (consistency with utility maximization) of GPTs directly (see also Jiang et al. (2025) for a survey) and Mazeika et al. (2025) elicits the inherent value functions of various LLMs. We provide a broader discussion of related research in Appendix A. In general, existing work tends to focus narrowly on a single theory, relies on limited model coverage, overlooks the effects of choice representation and output format, or lacks matched human comparisons. Our work complements these limitations.
@@ 2 Study Design
We study LLM decision-making under uncertainty. We consider a suite of 20 LLMs (detailed in the Supplement, Appendix B.1) designed to (i) cover widely used frontier models and (ii) enable controlled comparisons using open-weight models spanning different sizes and training stages. LLMs are queried through a unified interface with a minimal instruction-style template; details are in the Supplement (Appendix B.3). In addition, we analyze two references. The first is human subject behavior (see Section 2.2). The second is a risk-neutral rational agent (economicus) that always selects the option with a higher expected payoff.
@@ 2.1 Choices Among Prospects
We use three base prospect pairs—i.e., three two-option payoff distributions spanning distinct outcome scales—as a small, structured testbed for our manipulations. The pairs are intentionally close to (but not duplicates of) classic behavioral-economics stimuli, reducing the chance of LLM memorization (Kahneman and Tversky, 2013; Peterson et al., 2021); see Appendix B.4 for details.
Description-History Gap. Motivated by the description–experience (DE) gap in human behavior, we consider two ways of presenting prospects. The first is an explicit description, in which each prospect specifies a sequence of payoff–probability pairs; we refer to these as explicit prospects. The second consists of implicit prospects. While prior literature studies subjects who acquire experience by repeatedly selecting between prospects and observing stochastic payoff outcomes, we avoid the confound introduced by LLMs engaging in dynamic decision-making. Instead, we simulate experience by presenting prospects as simulated sequences of payoffs (histories) generated through repeated selection. We refer to the behavior discrepancy between these two representations as the description–history (DH) gap. We consider samples of 20 and 100 payoffs; results reported below aggregate across these samples, with a breakdown by sample size provided in Appendix C.3.
@@ 2.3 Interpretable Models of Behavior
To complement our analysis of the raw behavioral data, we make use of an interpretable model of (LLM or human) behavior to obtain deeper insights into behavior and effective risk preferences. The model—presented in its full form in the Supplement (Appendix B.2.1)—involves four nonnegative parameters: $\sigma$, which captures risk preference, $\lambda$ capturing loss aversion, $\gamma$, which captures probability weighting, and $\beta$, which corresponds to decisiveness. In particular:
@@
  • Higher $\lambda$ means greater sensitivity to loss (relative to gain) framing, $\lambda\rightarrow 0$ implies high sensitivity to gain, and $\lambda=1$ is insensitive to framing.
@@
  • Higher $\beta$ indicates greater determinism in behavior, while $\beta\rightarrow 0$ implies essentially uniformly random choice between two prospects.
To avoid pathologies, we bound all parameters in $[0.01,1000]$. An economicus then roughly corresponds to $\sigma=\lambda=\gamma=1$ and $\beta=1000$.
@@ 2.4 Quantifying Behavior
Querying and parsing. For each model and instantiated condition, we estimate a response distribution through repeated querying of ten times under identical inputs. We use temperature $T=1$ and a maximum generation length of 1024 tokens. If an LLM output does not contain a valid prospect choice, we retain the raw output and mark the trial as invalid. Choice rates are computed over valid trials.
@@
However, when MSE is non-zero, it may fail to capture an important property that decision patterns “go together”, i.e., the extent to which an increase or decrease in $p_{m}$ across difference contexts is accompanied by a similar change in $p_{m^{\prime}}$. To capture this, we treat $p_{m}(x)$ as a random variable with context $x$ viewed as the associated outcome. We can then measure similarity between a pair of models $m$ and $m^{\prime}$ using Pearson correlation between responses over all contexts $p_{m}(X)$ and $p_{m^{\prime}}(X)$, $\mathit{Cor}(p_{m}(X),p_{m^{\prime}}(X))$. If $m$ is an interpretable parametric model fit to data, we evaluate both MSE and correlation on a held out (test) dataset that is distinct from the data on which the parameters of $m$ where fit. Since the ability to capture the trends of behavior across contexts $x$ is essential to our analysis, we use correlation as the primary measure of both (a) similarity of pairs of models (e.g., LLMs), and (b) goodness of fit of a parametric model. Nevertheless, we observe that both MSE and correlation typically lead to the same goodness-of-fit conclusions.
Decisiveness and Consistency. In addition to model similarity and goodness of fit, two other behavioral trait we explore are decisiveness, which captures the entropy in the decisions, and consistency, of which we consider three forms. The first is order consistency, capturing the degree to which decisions tend to differ as a function of the order in which the prospects are presented. The second is prompt consistency, which measures the impact of explanation types on the nature of decisions. Finally, we consider frame consistency, which captures the impact of the gain vs. loss framing of otherwise identical prospects on the decision.
@@ 3 Behavior Patterns and Convergence
An immediate observation from the figure is the emergence of two LLM clusters: one that includes the (frontier) reasoning models GPT-5.1, DeepSeek-R1, and Gemini-2.5-Pro, and the other including their conversational counterparts (GPT-4.1, DeepSeek-Chat, and Gemini-2.5-Flash), as well as Claude-4.5-Haiku. In particular, correlations among pairs of models within each cluster are >0.8>0.8, and in most cases $\approx 0.9$ or higher. In contrast, correlations among pairs of models across clusters tend to be considerably lower—typically $\approx 0.6$ or below. This appears to be an example of convergence of frontier LLMs (Smith et al., 2025; Zhou et al., 2025; Mazeika et al., 2025), albeit across two distinct lines that reflect their target use. Henceforth, we thus distinguish between reasoning models (RMs) and conversational models (CM) in terms of their respective decision-making behavior under uncertainty.
@@
On the other hand, RMs are highly similar to the economicus reference, with correlations ranging between 0.86 and 1. In contrast, CMs exhibit low similarity to economicus, with correlations below 0.5. We also note that human behavior is essentially uncorrelated with economicus, with correlation only 0.14. This is largely a result of our selection of prospects that most emphasize the deviation of human behavior from expected utility maximization.
A split into RMs and CMs is also observed in open models; see the full heatmap in the Supplement (Appendix C.1). To simplify presentation, in Figure 2, we consider each model on a human-economicus scale, with the two references providing the two dimensions, and each 2D point reflecting model correlation to each of these We refer to this as the HE representation of LLMs, as well as the human subjects (who have a non-zero correlation to the economicus).
@@
The RM and CM clusters emerge here as well, with open models mapping predictably to each. For example, Qwen2.5-7B-Instruct is a CM model, while its mathematical reasoning variant is clearly an RM model (being far more economicus-like and somewhat less human-like than the instruct version). We see the same pattern with Qwen3-30B as well as with Olmo-3-7B (although here, the “Think” variant is slightly more human-like). In general, we observe that RMs tend to be far more economicus-like, and (usually) slightly less human-like compared to CMs.
@@ 4 Description-History Gap in LLMs
The results, presented as the HE representation of each model as well as the human subjects, are provided in Figure 4 for the frontier LLM models and in Figure 5 for open LLMs. Consider first the frontier LLMs, for which the observations are more crisp (Figure 4). An immediate and rather striking observation is that all models—whether RMs (x’s) or CMs (circles)—a change from explicit to implicit prospects increases the similarity to human behavior, and lowers similarity to economicus. An even more striking observation is the difference in how much the change to implicit prospects impacts RMs and CMs. In the case of the former, the impact tends to be moderate. For example, neither GPT-5.1 nor DeepSeek-R1 become especially less economicus-like when deciding from experience, though the change is more notable with Gemini Pro. For CMs, however, the change is dramatic: they become significantly more human-like, and significantly less economicus-like. Indeed, while RMs remain similar to one another in either setting, CMs occupy a relatively broad HE representation range under explicit prospects, but cluster closely with implicit prospects. For reference, we also show the change for the human subjects, who, somewhat surprisingly, become more economicus-like when prospects are presented implicitly; this likely accounts for the increased LLM-to-human correlation for RMs with implicit prospects.
@@ 5 Explaining Decisions
We investigate this by prompting LLMs for two kinds of explanations: a short (one-sentence) explanation and a full mathematical reasoning trace. Figure 6 visualizes the impact of these two explanation modes. What emerges is a surprising pattern, captured by the sequence of arrows starting at “no explanation” (red), then pointing to “short explanation” (blue), and then to “math explanation” (green). In particular, short explanations increase the degree of rationality (i.e., correlation with economicus) of most models, while asking for a longer rationale often reduces decision rationality in comparison with the former. Notably, this pattern is largely consistent for both frontier and open CMs, with Claude-4.5-Haiku somewhat of an outlier, whose both types of explanations resulting in behavior that is less like economicus compared to no explanation provided. Trends in RM behavior across explanation modes are more heterogeneous: More than half of the models (GPT-5.1, DeepSeek-R1, Olmo3-7B-think, and Qwen2.5-Math-7B-Instruct) exhibit strong stability; Gemini-2.5-Pro follows a pattern similar to the CMs; In contrast, Qwen3-30B-Thinking and Olmo3-32B-Think show significant changes when prompted to provide a mathematical explanation, while no differences are observed between the no-explanation and brief-explanation modes. One possible explanation is that, because CMs are not trained for mathematical reasoning, prompting them to provide mathematical explanations may only induce verbosity but no improved rationality.

### v2

@@ Abstract
The use of large language models either as decision support systems, or in agentic workflows, is rapidly transforming the digital ecosystem. However, the understanding of LLM decision-making under uncertainty remains limited. We study LLM risky choices along two dimensions: (1) prospect representation (based on an explicit representation or outcome history) and (2) decision rationale (explanation). Our study, which involves 20 frontier and open LLMs, is complemented by a matched human subjects experiment, which provides one reference point, while an expected payoff maximizing rational agent model provides another. We find that LLMs cluster into two categories: reasoning models (RMs) and conversational models (CMs). RMs tend towards rational behavior, are insensitive to the order of prospects, gain/loss framing, and explanations, and behave similarly whether prospects are explicit or presented via a history of outcomes. CMs are significantly less rational, slightly more human-like, sensitive to prospect ordering, framing, and explanation, and exhibit a large description-history gap. Paired comparisons of open LLMs suggest that a key factor differentiating RMs and CMs is training for mathematical reasoning.
@@ 1 Introduction
However, LLMs are neither explicitly designed to maximize expected payoff, nor necessarily behave like humans. Rather, their competence is an emergent property of scale and training paradigms including next-word prediction, instruction fine-tuning, human preference alignment, and mathematical reasoning (Du et al., 2024; Guo et al., 2025; Ouyang et al., 2022). The net effect of this soup of training methods on LLM economic decision-making under uncertainty remains unclear. In response, a literature has begun to emerge that aims to investigate LLM decision-making, particularly under uncertainty Binz and Schulz (2023); Coda-Forno et al. (2024); Jia et al. (2024); Ross et al. (2024); Payne (2025); Wang et al. (2025). However, there remain conflicting accounts, for example, of the extent LLMs are rational or align with prospect theory and human behavior (Chen et al., 2023; Horton, 2023; Jia et al., 2024; Ross et al., 2024; Wang et al., 2025; Payne, 2025). More fundamentally, given the broad variety of training paradigms, goals, and architectures, it is unclear whether a straightforward account of LLM economic decision making is even possible.
@@
We consider 20 frontier and open LLMs, along with human subjects’ choices over the same prospects and treatments. LLMs are analyzed vis-a-vis two references: 1) the human subjects pool, and 2) the idealized economicus rational agent. Within the scope of our experiments, our key findings are:
@@
    All models exhibit a description-history (DH) gap, an analogue of the decision-experience (DE) gap in which agents only see a history of past outcomes. This gap is considerable for CMs, but modest for RMs.
@@
Related Work. LLMs are increasingly studied as objects of behavioral analysis Binz and Schulz (2023); Coda-Forno et al. (2024); Dillion et al. (2023); Hagendorff et al. (2023); Hayes et al. (2024); Horton (2023); Ivanova (2023). Jia et al. (2024) analyze LLM decisions by eliciting choices and fitting prospect-theoretic models; subsequent work extends this by examining epistemic markers Wang et al. (2025), application contexts Payne (2025), and persona matching  Liu et al. (2025). Horowitz and Plonsky (2025) study a description-experience gap in LLMs through repeated choice, unlike the passive sampling used in our setting. Apart from the bounded rationality focus, Chen et al. (2023) instead focuses on rationality (consistency with utility maximization) of GPTs directly (see also Jiang et al. (2025) for a survey) and Mazeika et al. (2025) elicits the inherent value functions of various LLMs, while Coda-Forno et al. (2024) provide a benchmarking suite for LLM decision making under risk. We provide a broader discussion of related research in Appendix A. In general, existing work tends to focus narrowly on a single theory, relies on limited model coverage, overlooks the effects of choice representation and output format, or lacks matched human comparisons. Our work aims to address some of these limitations.
@@ 2 Study Design
We investigate how large language models (LLMs) make decisions when faced with uncertainty. To do this comprehensively, we examine 20 different LLMs (detailed in the Supplement, Appendix B.1). Our selection covers two important goals: first, we include widely used frontier models; second, we use open-weight models of varying sizes and training stages to enable controlled comparisons. We query all LLMs through a unified interface using a minimal instruction template (see Supplement, Appendix B.3 for our prompts and an additional prompt sensitivity analysis). To measure behavior, we compare LLM responses against two references: human subject behavior and economicus, a risk-neutral expected payoff maximizer.
@@ 2.1 Choices Among Prospects
Our study uses three base prospect pairs as a structured testbed. These pairs span different outcome scales and are intentionally similar to (but not duplicates of) classic behavioral-economics stimuli (Kahneman and Tversky, 2013; Peterson et al., 2021). This design reduces the risk of LLM memorization. See Appendix B.4 for details.
Description-History Gap. Human decision-making differs depending on how information is presented. Research in the description–experience (DE) gap shows that people decide differently when given explicit information versus when they learn through experience. We adapt this insight to study LLMs. Rather than having LLMs interact dynamically with an environment (which would introduce confounds), we present prospects in two ways:
  • Explicit Prospects. Each prospect specifies exact payoff–probability pairs. For example: “70% chance of \$100, 30% chance of \$0.”
  • Implicit Prospects. We present prospects as simulated histories – sequences of payoffs that would result from repeated selection. For instance, a simulated history might show 15 instances of \$100 and 5 instances of \$0 from 20 trials, representing the same underlying distribution.
We call the behavioral difference between these presentations the description–history (DH) gap. We test with simulated histories of 20 and 100 payoffs, with results aggregated across sample sizes (see Appendix C.3 for breakdowns by size).
@@ 2.3 Interpretable Models of Behavior
To complement our analysis of the raw behavioral data, we make use of interpretable models of (LLM or human) behavior to obtain deeper insights into behavior and effective risk preferences. We consider two parameterizations of prospect theory, both with four free parameters (presented in full in Appendix B.2.1). The first is a standard formulation with parameters $\sigma$ (risk preference), $\lambda$ (loss aversion), $\gamma$ (probability weighting), and $\beta$ (decisiveness). However, because $\lambda$ and $\beta$ are not jointly identifiable in pure-loss prospects, we also consider an alternative specification as our primary model (a special case of the generalized power value model Peterson et al. (2021)), replacing $\lambda$ and $\beta$ with $\beta_{\text{gain}}$ and $\beta_{\text{loss}}$ for the gain and loss frames, respectively, while retaining $\sigma$ and $\gamma$. In these models:
@@
  • Higher $\beta_{\text{gain}}$ (and $\beta_{\text{loss}}$) indicates greater determinism in the gain (and loss) domain, while $\beta\rightarrow 0$ implies essentially uniformly random choice. The ratio $\beta_{\text{loss}}/\beta_{\text{gain}}$ serves as a proxy for loss aversion: a ratio greater than 1 indicates higher sensitivity to utility differences in losses than in gains.
To avoid pathologies, we bound all parameters in $[0.01,1000]$. An economicus then roughly corresponds to $\sigma=\gamma=1$ and $\beta_{\text{gain}}=\beta_{\text{loss}}=1000$.
@@ 2.4 Quantifying Behavior
Querying and parsing. For each model and instantiated condition, we estimate a response distribution through repeated querying of 10 times under identical inputs. We use temperature $T=1$ and a maximum generation length of 1024 tokens. We found that increasing the number of samples to 50 has minimal qualitative impact on the results, while being significantly more costly. If an LLM output does not contain a valid prospect choice, we retain the raw output and mark the trial as invalid. Choice rates are computed over valid trials.
@@
However, when MSE is non-zero, it may fail to capture an important property that decision patterns “go together”, i.e., the extent to which an increase or decrease in $p_{m}$ across different contexts is accompanied by a similar change in $p_{m^{\prime}}$. To capture this, we treat $p_{m}(x)$ as a random variable with context $x$ viewed as the associated outcome. We can then measure similarity between a pair of models $m$ and $m^{\prime}$ using Pearson correlation between responses over all contexts $p_{m}(X)$ and $p_{m^{\prime}}(X)$, $\mathit{Cor}(p_{m}(X),p_{m^{\prime}}(X))$. If $m$ is an interpretable parametric model fit to data, we evaluate both MSE and correlation on a held out (test) dataset that is distinct from the data on which the parameters of $m$ were fit. Since the ability to capture the trends of behavior across contexts $x$ is essential to our analysis, we use correlation as the primary measure of both (a) similarity of pairs of models (e.g., LLMs), and (b) goodness of fit of a parametric model. Nevertheless, we observe that both MSE and correlation typically lead to the same goodness-of-fit conclusions.
Decisiveness and Consistency. In addition to model similarity and goodness of fit, two other behavioral traits we explore are decisiveness, which captures the entropy in the decisions, and consistency, of which we consider three forms. The first is order consistency, capturing the degree to which decisions tend to differ as a function of the order in which the prospects are presented. The second is prompt consistency, which measures the impact of explanation types on the nature of decisions. Finally, we consider frame consistency, which captures the impact of the gain vs. loss framing of otherwise identical prospects on the decision.
@@ 3 Behavior Patterns and Convergence
An immediate observation from the figure is the emergence of two LLM clusters: one that includes the (frontier) reasoning models GPT-5.1, DeepSeek-R1, and Gemini-2.5-Pro, and the other including their conversational counterparts (GPT-4.1, DeepSeek-Chat, and Gemini-2.5-Flash), as well as Claude-4.5-Haiku. In particular, correlations among pairs of models within each cluster are >0.8>0.8, and in most cases $\approx 0.9$ or higher. In contrast, correlations among pairs of models across clusters tend to be considerably lower—typically $\approx 0.6$ or below. This appears to be an example of convergence of frontier LLMs (Smith et al., 2025; Zhou et al., 2025; Mazeika et al., 2025), albeit across two distinct lines that reflect their target use. Henceforth, we thus distinguish between reasoning models (RMs) and conversational models (CM) in terms of their respective decision-making behavior under uncertainty. This split into RMs and CMs is also observed in open models; see the full heatmap in the Supplement (Appendix C.1).
@@
On the other hand, RMs are quite similar to the economicus reference, with correlations ranging from 0.86 to 1. In contrast, CMs exhibit low similarity to economicus, with correlations below 0.5. We also note that human behavior is essentially uncorrelated with economicus, with correlation only 0.14. This is likely a consequence of selecting prospects that most emphasize the deviation of human behavior from expected utility maximization.
To confirm that these structural differences are robust to sampling noise, we conducted an empirical bootstrap analysis: for each model and prospect pair, we resampled 1,000 synthetic datasets from a Binomial distribution parameterized by the observed choice frequency, and recomputed all behavioral metrics. The RM–CM clustering, as well as the deviation of human behavior from both, hold with $p<0.001$ across all comparisons.
To simplify presentation, in Figure 2, we introduce the HE representation, plotting each model on a two-dimensional space anchored by human and economicus correlations (note humans have a non-zero correlation to the economicus). This visualization provides a compact, interpretable framework for understanding LLM behavior relative to two behavioral extremes. The RM and CM clusters emerge here as well, with open models mapping predictably to each. For example, Qwen2.5-7B-Instruct is a CM model, while its mathematical reasoning variant is clearly an RM model (being far more economicus-like and somewhat less human-like than the instruct version). We see the same pattern with Qwen3-30B as well as with Olmo-3-7B (although here, the “Think” variant is slightly more human-like). In general, we observe that RMs tend to be far more economicus-like, and (usually) slightly less human-like compared to CMs.
@@ 4 Description-History Gap in LLMs
The results, presented as the HE representation of each model as well as the human subjects, are provided in Figure 4 for the frontier LLM models and in Figure 5 for open LLMs. Consider first the frontier LLMs, for which the observations are more crisp (Figure 4). An immediate and rather striking observation is that for all models—whether RMs (x’s) or CMs (circles)—the change from explicit to implicit prospects increases the similarity to human behavior, and lowers similarity to economicus. An even more striking observation is the difference in how much the change to implicit prospects impacts RMs and CMs. In the case of the former, the impact tends to be moderate. For example, neither GPT-5.1 nor DeepSeek-R1 become especially less economicus-like when deciding from experience, though the change is more notable with Gemini Pro. For CMs, however, the change is dramatic: they become significantly more human-like, and significantly less economicus-like. Indeed, while RMs remain similar to one another in either setting, CMs occupy a relatively broad HE representation range under explicit prospects, but cluster closely with implicit prospects. For reference, we also show the change for the human subjects, who, somewhat surprisingly, become more economicus-like when prospects are presented implicitly; this likely accounts for the increased LLM-to-human correlation for RMs with implicit prospects.
@@
We note that all the models we tested have very long context windows. For example, DeepSeek-Chat has a 128K context limit. Even a 100-sample history represents only a small fraction of this available context, making it unlikely that the observed gap stems from context length limitations. This is further supported by results in the Appendix C showing that many LLMs exhibit more rational behavior with longer (100-sample) histories than with shorter (20-sample) histories, providing strong evidence that context constraints are not the underlying cause of the DH gap.
@@ 5 Explaining Decisions
We investigate this by prompting LLMs for two kinds of explanations: a short (one-sentence) explanation and a full mathematical reasoning trace. Figure 6 visualizes the impact of these two explanation modes. What emerges is a surprising pattern, captured by the sequence of arrows starting at “no explanation” (red), then pointing to “short explanation” (blue), and then to “math explanation” (green). In particular, short explanations increase the degree of rationality (i.e., correlation with economicus) of most models, while asking for a longer rationale often reduces decision rationality in comparison with the former. Notably, this pattern is largely consistent for both frontier and open CMs, with Claude-4.5-Haiku somewhat of an outlier, whose both types of explanations result in behavior that is less like economicus compared to no explanation provided. Trends in RM behavior across explanation modes are more heterogeneous: More than half of the models (GPT-5.1, DeepSeek-R1, Olmo3-7B-think, and Qwen2.5-Math-7B-Instruct) exhibit strong stability; Gemini-2.5-Pro follows a pattern similar to the CMs; In contrast, Qwen3-30B-Thinking and Olmo3-32B-Think show significant changes when prompted to provide a mathematical explanation, while no differences are observed between the no-explanation and brief-explanation modes. One possible explanation is that, because CMs are not trained for mathematical reasoning, prompting them to provide mathematical explanations may only induce verbosity but no improved rationality.